Write all the factors of the following numbers: 
27 ??????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

27 = 1 × 27 27 = 3 × 9 27 = 9 x 3
Stop here since 3 and 9 have occurred earlier
Hence, the factors of 27 are 1, 3, 9 and 27

Thanks